HELIUM NANODROPLET ISOLATION SPECTROSCOPY OF NO_{2} AND VAN DER WAALS COMPLEXES

\underline{ROBERT R. FEHNEL}, KEVIN K. LEHMANN, Department of Chemistry, University of Virginia, Charlottesville VA, 22904-4319.

We will report the infrared laser spectra of NO_{2}, NO_{2}-N_{2}, NO_{2} - H_{2}O and NO_{2} - O_{2} complexes using Helium Nanodroplet Isolation (HENDI) Spectroscopy. These molecules were excited using an ARGOS cw PPLN OPO (Aculight) in the 3.2 - 3.9 μm spectral window and we excited the ν_{1} + ν_{3} transition of NO_{2}.